# A continuum mechanics approach for the deformation of non-Euclidean origami generated by piecewise constant nematic director fields

### v1

@@ 4.2.3 Admissible domain to avoid self-intersection
In our previous analysis, we choose $\gamma_{2}$ as the reference folding angle, and the folding angle functions ensure the continuity of the deformation ${\bf y}(\tilde{{\bf x}})$. However, this analysis allows unphysical self-intersections, and one panel may be folded to pass through its adjacent panels. Thus, we need to determine the admissible domain of the reference folding angle. The domain ${\cal I}^{+}\cup{\cal I}^{-}$ is obtained only from the existence of $\gamma_{4}$. Therefore, the admissible domain is a subset of ${\cal I}^{+}\cup{\cal I}^{-}$ such that no self-intersection occurs at any crease.
Notice that the folding angle jumps between $-\pi$ and $\pi$ when the panel intersects with its adjacent panel. Thus, we perform the following check
@@
to determine the reference folding angle $\bar{\theta}$ at which the folding angles $\gamma_{i}$ reaches ±$\pm\pi$. Let ${\cal J}^{+}\subset{\cal I}^{+}=[\theta_{\min},\theta_{\max}]$ be the set of points at which one of the folding angles jumps, say,
@@
as the upper bound of the physically relevant reference folding angle (i.e., no self-intersection) on ${\cal I}^{+}$. Similarly, the lower bound $\bar{\theta}_{\min}$ on ${\cal I}^{-}=[-\theta_{\max},-\theta_{\min}]$ is defined as
@@
${\cal J}^{-}:=\{\theta\in{\cal I}^{-}:|\gamma_{i}(\theta^{+})-\gamma_{i}(\theta^{-})|=2\pi,~{}i=1,2,3~{}\text{or}~{}4\}.$ (48)
@@
¯$\bar{{\cal I}}^{+}=[\theta_{\min},\bar{\theta}_{\max}],\quad\bar{{\cal I}}^{-}=[\bar{\theta}_{\min},-\theta_{\min}].$ (49)
In reality, the domain can be obtained numerically by simply performing the above check, resulting in the folding angle curves in Fig. 7 that do not exceed ±$\pm\pi$.
@@
$\gamma_{1}=\bar{\gamma}_{1}^{\tilde{\sigma}(\theta)}(\theta,\sigma),~{}\gamma_{2}=\theta,~{}\gamma_{3}=\bar{\gamma}_{3}^{\sigma}(\theta),~{}\gamma_{4}=\bar{\gamma}_{4}^{\sigma}(\theta),~{}\sigma\in\{\pm 1\},~{}\theta\in\bar{{\cal I}}^{-}\cup\bar{{\cal I}}^{+}.$ (50)
@@
Remark 1: Symmetry relation. By checking the folding angle functions, we claim that if $(\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4})$ is on the branch $\sigma=+$, then $(-\gamma_{1},-\gamma_{2},-\gamma_{3},-\gamma_{4})$ is on the other branch $\sigma=-$, as shown in Fig. 7. These two configurations are mirror related.
@@
Remark 3: Connection to director fields. In the four-fold case, we do not have a direct relation between the director pattern on the reference domain and the folding angles on the folded configuration, as in the three-fold case. However, the numerical approach is similar. First, a topology condition is needed ($({\bf t}_{i}\times{\bf t}_{i+1})\cdot{\bf e}_{3}=+1$) for the initial director pattern. Then we can calculate the deformed sector angles by $\tilde{\alpha}_{i}=\arccos[({\bf U}_{{\bf n}_{i}}{\bf t}_{i}\cdot{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1})/(|{\bf U}_{{\bf n}_{i}}{\bf t}_{i}||{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1}|)]$. Substituting these deformed sector angles into the folding angle functions yields the kinematics of the four-fold origami.
@@ 5.1 Quadrilateral director patterns
Figure 8: (a) The director on the $(3,3)$ panel (in the red frame) is determined by the other directors and the folding angles $\gamma_{3},\gamma_{4}$. (b) A marching algorithm: the director pattern within the dashed red frame is determined by the boundary directors and the reference folding angle, following (a).
@@
To design a quadrilateral pattern, we prescribe the directors and the corresponding tangents at the left and bottom boundaries (panels with indices $(i,1),(i,2),(1,j),(2,j)$, i.e. the domain outside the red dashed lines in Fig. 8(b)). By entering a reference folding angle $\theta$ and applying Eq. (54), the pattern within the red dashed domain is determined. Once all directors are determined, we fix the tangent branches and compute the tangents. The coordinates of the vertices are then computed by these tangents. The difficulty we might have is that the process may break the topology of a quadrilateral mesh. Thus, a more realistic way is to start with a symmetric pattern and then perturb the boundary directors. Figure 8(b) is an example of irregular quadrilateral patterns. We start with a symmetric pattern with directors along ${\bf R}_{{\bf e}_{3}}(\pi/4){\bf e}_{1}$ and ${\bf R}_{{\bf e}_{3}}(3\pi/4){\bf e}_{1}$. As shown in Fig. 8(b), the symmetric pattern consists of Type A nodes and Type B nodes, which will deform into pyramids and saddles, respectively. We then perturb the directors at the boundary vertices and compute the entire pattern using Eq. (54). During the marching process, the folding angle branches are fixed by $\sigma=+1$ in Eq. (50), meaning the opposite folding angles have the same sign. The resulting quadrilateral pattern is irregular in both the reference and actuated configurations, and the distribution of Gaussian curvature is non-symmetric. In the deformed configuration, we observe alternating positive and negative Gaussian curvature at the four-fold vertices.
@@ 5.2 Director patterns combining three-fold and four-fold intersections
Figure 9: (a) Notation of the unit cell. (b) Notation of the four-fold intersection in the pattern. (c) The designed pattern composed of three-fold and four-fold intersections. The pattern is compatible in both the reference and actuated configurations.
@@
for any $\theta_{i}$ satisfying the topological condition ${\bf t}_{i}\times{\bf t}_{i+1}\cdot{\bf e}_{3}=1$, where ${\bf R}_{i}$ is a rotation tensor with rotation axis $\tilde{{\bf t}}_{i}$ for the intermediate configuration. The numerical result given by the global minimization NMaximize in Mathematica shows that $error=0$ in the domain $0<\theta_{1}<\theta_{2}<\theta_{3}<\theta_{4}<2\pi$, indicating that the four-fold vertex is compatible in our setting. Since compatibility holds for arbitrary four-fold nodes, we can successively stack different unit cells along the i- and j-directions to construct both the reference pattern and the actuated configuration, as shown in Fig. 9(c). In the actuated state, the configuration exhibits positive Gaussian curvature at the three-fold vertices and negative Gaussian curvature at the four-fold vertices. Notably, the design space for the director fields is quite large; the primary constraint is the preservation of the pattern’s topological structure.
@@ 6 Discussion
To tackle this problem, we first analyze the fundamental condition of metric compatibility between adjacent director fields. We then develop a rigorous continuum mechanics framework to study the kinematics of isolated three-fold and four-fold vertices. Building on this vertex-level understanding, we construct two classes of large-scale director patterns that are compatible both before and after actuation. In other words, the directors satisfy metric compatibility at each crease, and the resulting global deformation is geometrically compatible. Although the general compatibility problem remains open, our two proposed designs demonstrate broad and flexible design spaces. The first is a quadrilateral director pattern with free directors on the two boundaries. Using a marching algorithm and a rigidity theorem, the director information propagates to interior nodes, producing a compatible actuated configuration characterized by alternating positive and negative Gaussian curvature at the vertices. The second design combines three-fold and four-fold intersections, leveraging a folding angle theorem that shows the folding angle at a three-fold node depends solely on its adjacent directors. In this configuration, positive Gaussian curvature arises at the three-fold vertices and negative curvature at the four-fold vertices after actuation.
@@ Acknowledgments
This work is based on discussions with Prof. Paul Plucinsky at USC, while we were working together at the University of Minnesota. F.F. acknowledges the financial support from the National Natural Science Foundation of China (Grant No. 12472061).

### v2

@@ 5.2.3 Admissible domain to avoid self-intersection
[…]
Notice that the folding angle jumps between $\pi$ and $\pi$ when the panel intersects with its adjacent panel. Thus, we perform the following check
@@
to determine the reference folding angle $\bar{\theta}$ at which the folding angles $\gamma_{i}$ reaches $\pm\pi$. Let ${\cal J}^{+}\subset{\cal I}^{+}=[\theta_{\min},\theta_{\max}]$ be the set of points at which one of the folding angles jumps, say,
@@
[…]
@@
${\cal J}^{-}:=\{\bar{\theta}\in{\cal I}^{-}:|\gamma_{i}(\bar{\theta}^{+})-\gamma_{i}(\bar{\theta}^{-})|=2\pi,~i=1,2,3~\text{or}~4\}.$ (52)
@@
$\bar{{\cal I}}^{+}=[\theta_{\min},\bar{\theta}_{\max}],\quad\bar{{\cal I}}^{-}=[\bar{\theta}_{\min},-\theta_{\min}].$ (53)
In reality, the domain can be obtained numerically by simply performing the above check, resulting in the folding angle curves in Fig. 9 that do not exceed $\pm\pi$.
@@
$\gamma_{1}=\bar{\gamma}_{1}^{\tilde{\delta}(\theta)}(\theta,\delta),~\gamma_{2}=\theta,~\gamma_{3}=\bar{\gamma}_{3}^{\delta}(\theta),~\gamma_{4}=\bar{\gamma}_{4}^{\delta}(\theta),~\delta\in\{\pm 1\},~\theta\in\bar{{\cal I}}^{-}\cup\bar{{\cal I}}^{+}.$ (54)
@@
Remark 1: Symmetry relation. By checking the folding angle functions, we claim that if $(\gamma_{1},\gamma_{2},\gamma_{3},\gamma_{4})$ is on the branch $\delta=+$, then $(-\gamma_{1},-\gamma_{2},-\gamma_{3},-\gamma_{4})$ is on the other branch $\delta=$-, as shown in Fig. 9. These two configurations are mirror related.
@@
5.3 Connection to director fields
As discussed in Section 3, for the four-fold intersection, we only need to consider the degenerate case satisfying $\alpha_{1}+\alpha_{3}=\pi$ (or equivalently $\alpha_{2}+\alpha_{4}=\pi$). The example in Fig. 10 also satisfies this condition. Assuming ${\bf n}_{1}={\bf R}_{{\bf e}_{3}}(\theta){\bf e}_{1}$, we compute the director at each panel, and then calculate the deformed sector angles by $\tilde{\alpha}_{i}=\arccos[({\bf U}_{{\bf n}_{i}}{\bf t}_{i}\cdot{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1})/(|{\bf U}_{{\bf n}_{i}}{\bf t}_{i}||{\bf U}_{{\bf n}_{i}}{\bf t}_{i+1}|)]$. Substituting these deformed sector angles into the folding angle functions yields the kinematics of the four-fold origami directly from the director pattern.
Suppose the reference sector angles are given by $(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(\alpha_{1},\alpha_{2},\pi-\alpha_{1},\pi-\alpha_{2})$, ${\bf n}_{1}={\bf R}_{{\bf e}_{3}}(\theta){\bf e}_{1}$ and ${\bf t}_{1}={\bf e}_{1}$. By using ${\bf n}_{i+1}={\bf n}_{i}-2({\bf n}_{i}\cdot{\bf t}_{i+1}){\bf t}_{i+1}$, we compute the directors ${\bf n}_{i}={\bf R}_{{\bf e}_{3}}(\theta_{i}){\bf e}_{1}$ as
$\theta_{1}=\theta,~\theta_{2}=2\alpha_{1}+2\alpha_{2}-\theta+\pi,~\theta_{3}=\theta-2\beta_{1},~\theta_{4}=2\beta_{1}-\theta+\pi.$ (56)
We then substitute the directors and the crease tangents to obtain the deformed sector angles $\tilde{\alpha}_{i}$. A further calculation of Eq. (36) yields (numerically validated)
$\cos(\gamma_{4})$ $=$ $\frac{\cos(\gamma_{2})\!\left[1+r^{2}+(-1+r^{2})\cos(4\alpha_{1}-2\theta)\right]-2(-1+r^{2})\sin(\alpha_{1}-\alpha_{2})\sin(3\alpha_{1}+\alpha_{2}-2\theta)}{1+r^{2}+(-1+r^{2})\cos\!\big[2(\alpha_{1}+\alpha_{2}-\theta)\big]}$ (57)
$:=$ $\tilde{F}_{4}(\gamma_{2};\boldsymbol{\alpha},\theta)$
where $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4})=(\alpha_{1},\alpha_{2},\pi-\alpha_{1},\pi-\alpha_{2})$ and $r=\lambda_{\parallel}/\lambda_{\perp}$. Equation (57) gives the relationship between the opposite folding angles $\gamma_{4}$ and $\gamma_{2}$ in terms of the directors and creases (determined by $\theta$ and $\alpha_{i}$). Unfortunately, there is no such neat formula for the adjacent folding angles. But we can always obtain the deformed sector angles $\tilde{\alpha}_{i}$ in terms of $\theta$ and $\alpha_{i}$, and then compute the kinematics, because all the formulae including the folding angle functions (44) and (36) are explicit. The explicit forms of these functions provide the foundation for efficient numerical algorithms used to pattern the complex director fields in the next section.
@@ 6.1 Quadrilateral director patterns
Figure 10: (a) The director on the $(3,3)$ panel (in the red frame) is determined by the other directors and the folding angles $\gamma_{3},\gamma_{4}$. (b) A marching algorithm: the director pattern within the dashed red frame is determined by the boundary directors and the reference folding angle, following (a). (c) The orthogonal dual of (b). The patterns are actuated by $\lambda_{\parallel}=0.9,\lambda_{\perp}=1.1$. The blue and red spheres at the vertices represent negative and positive Gaussian curvature.
@@
To design a quadrilateral pattern, we prescribe the directors and the corresponding tangents at the left and bottom boundaries (panels with indices $(i,1),(i,2),(1,j),(2,j)$, i.e. the domain outside the red dashed lines in Fig. 10(b)). By entering a reference folding angle $\theta$ and applying Eq. (60), the pattern within the red dashed domain is determined. Once all directors are determined, we fix the tangent branches and compute the tangents. The coordinates of the vertices are then computed by these tangents. The difficulty we might have is that the process may break the topology of a quadrilateral mesh. Thus, a more realistic way is to start with a symmetric pattern and then perturb the boundary directors. Figure 10(b) is an example of irregular quadrilateral patterns. We start with a symmetric pattern with directors along ${\bf R}_{{\bf e}_{3}}(\pi/4){\bf e}_{1}$ and ${\bf R}_{{\bf e}_{3}}(3\pi/4){\bf e}_{1}$. As shown in Fig. 10(b), the symmetric pattern consists of Type A nodes and Type B nodes, which will deform into pyramids and saddles, respectively. We then perturb the directors at the boundary vertices and compute the entire pattern using Eq. (60). During the marching process, the folding angle branches are fixed by $\delta=+1$ in Eq. (54), meaning the opposite folding angles have the same sign. The resulting quadrilateral pattern is irregular in both the reference and actuated configurations, and the distribution of Gaussian curvature is non-symmetric. In the deformed configuration, we observe alternating positive and negative Gaussian curvature at the four-fold vertices. We also show the actuated configuration and Gaussian curvature of the orthogonal dual in Fig. 10(c).
We may wonder if the structures in Fig. 10 are foldable or not. If we fix $\lambda_{\parallel}=0.9,\lambda_{\perp}=1.1$, these actuated configurations are rigid, because numerically we find that a different reference folding angle leads to a different reference director pattern (except for periodic cases in Fig. 12). However, if we fix the reference folding angle and the reference director pattern, then gradually develop the stretches $\lambda_{\parallel}=\lambda,\lambda_{\perp}=\lambda^{-0.5}$ by decreasing $\lambda$ from 0.95 to 0.65, we observe a continuous compatible evolution (Fig. 11(a)). This observation relies on the fact that, if we assume the reference pattern is compatible and there exists a compatible deformed configuration for $\lambda_{\parallel,\perp}$, we can always find a compatible deformed configuration for other $\lambda_{\parallel,\perp}$. Though not rigorously proved, the conjecture is likely to be true based on numerical observations, as also shown in Figs. 12(d) and 13(d).
The above designs have $\delta=+1$ for the folding branch. Here we provide another example with a different folding branch $\delta$ to show the versatility of our design strategy. The two deformed configurations of the example in Fig. 11(b) share the same reference director pattern, actuation stretches $\lambda_{\parallel}=0.95,\lambda_{\perp}=1.05$, reference folding angle $\eta=0.8$, but have different folding branches $\delta$. The right deformed configuration has all $\delta=+1$, exhibiting shapes similar to those shown in Fig. 10. For the left deformed configuration, we switch the folding branch from $\delta=+1$ to $\delta=-1$ along a row of vertices, leading to sharp geometric changes. Since self-intersection easily occurs in this case, only a $3\times 3$ pattern is shown here. It should be noted that, unlike the rigidly and flat-foldable quadrilateral mesh origami in our previous work feng2020designs , we have no explicit formula for $\delta$ of the forth node as a function of the geometries and folding branches of the first three, by solving the compatibility equation in Fig. 10(a). However, it is numerically verified that we can always find a solution of $\delta$ for the fourth node as shown in Fig. 11(b), given the folding branches of the first three ($\delta=+1$ or $\delta=-1$).
Figure 11: (a) The evolution of the same director pattern by decreasing $\lambda$ from 0.95 to 0.65 where $\lambda_{\parallel}=\lambda,\lambda_{\perp}=\lambda^{-0.5}$.(b) The two deformed configurations share the same reference director pattern, actuation stretches $\lambda_{\parallel}=0.95,\lambda_{\perp}=1.05$, reference folding angle $\eta=0.8$, but have different folding branches $\delta$.
6.2 Periodic quadrilateral pattern and active metamaterial
Harnessing the unique properties of non-Euclidean origami, we may design active metamaterials xiao2020active ; dudek2025shape that are responsive to external stimuli. A convenient strategy is to use periodic patterns, as shown in Fig. 12. The reference domain is a periodic square pattern, with directors pointing along $\pm\pi/4$ at each panel (Fig. 12(a)). We first study the motion of a single pattern and then try to construct a metamaterial based on it in 3D. We also notice that the evolved structure is exactly the eggbox origami studied by Paulino in pratapa2019geometric .
Figure 12: (a) A periodic reference director pattern. (b) Metamaterials constructed by periodic non-Euclidean origami. The joined vertices remain connected during motion. (c)-(d) Two deformation modes regarding rigid folding and active evolution respectively. (c) Fix the stretches $\lambda_{\parallel},\lambda_{\perp}$ and change the reference folding angle $\eta$. (d) Fix the reference folding angle $\eta$ and change the stretches $\lambda_{\parallel}=\lambda,\lambda_{\perp}=\lambda^{-0.5}$ from $\lambda=0.95$ to $\lambda=0.5$.
The non-Euclidean origami has two modes of motion. First we assume the director pattern (each panel) is fully evolved by $\lambda_{\parallel}=0.9$ and $\lambda_{\perp}=1.1$ in Fig. 12(c), and then change the folding angle $\eta$ from $0.3$ to $1.5$. Owing to the symmetry, the origami is rigidly foldable, exhibiting an overall shape change from rectangle to square and to rectangle shape as shown in Fig. 12(c). The other mode is that we can fix the folding angle $\eta=1$ and evolve the stretches $\lambda_{\parallel}=\lambda,\lambda_{\perp}=\lambda^{-0.5}$ from $\lambda=0.95$ to $\lambda=0.5$. Unlike the first mode, the overall pattern evolves through large-square, small-square, and rectangular shapes. The second mode mimics the shape evolution during the phase transformation of liquid crystal elastomer sheets.
Since the origami pattern remains compatible and periodic during the two modes of motion, we could stack the actuated pattern in multilayers by joining corresponding nodes periodically (Fig. 12(b)). The stacked non-Euclidean origami forms an active metamaterial that can undergo shape change by folding it (mode 1) and by stimulating it (mode 2). The latter would equip our designed active metamaterial with better properties compared with classic passive metamaterial, since many physical properties related to shapes (e.g. band gaps, damping, Poisson’s ratio) can be modulated upon external stimuli in our case. A detailed study of the active metamaterial is out of the scope of this paper, and we reserve it for future work.
@@ 6.3 Director patterns combining three-fold and four-fold intersections
Figure 13: (a) Notation of the unit cell. (b) Notation of the four-fold intersection in the pattern. (c) The designed pattern composed of three-fold and four-fold intersections. The pattern is compatible in both the reference and actuated configurations. (d) The evolution of the structure by changing the stretches $\lambda_{\parallel}=\lambda,\lambda_{\perp}=\lambda^{-0.5}$ from $\lambda=1$ to $\lambda=0.5$.
@@
for any $\theta_{i}$ satisfying the topological condition ${\bf t}_{i}\times{\bf t}_{i+1}\cdot{\bf e}_{3}=1$, where ${\bf R}_{i}$ is a rotation tensor with rotation axis $\tilde{{\bf t}}_{i}$ for the intermediate configuration. The numerical result given by the global minimization NMaximize in Mathematica shows that $error=0$ in the domain $0<\theta_{1}<\theta_{2}<\theta_{3}<\theta_{4}<2\pi$, indicating that the four-fold vertex is compatible in our setting. Since compatibility holds for arbitrary four-fold nodes, we can successively stack different unit cells along the i- and j-directions to construct both the reference pattern and the actuated configuration, as shown in Fig. 13(c). In the actuated state, the configuration exhibits positive Gaussian curvature at the three-fold vertices and negative Gaussian curvature at the four-fold vertices. Notably, the design space for the director fields is quite large; the primary constraint is the preservation of the pattern’s topological structure. We also notice that the actuated configuration can be evolved in a compatible way by reducing $\lambda=1$ to $\lambda=0.5$ in the stretches $\lambda_{\parallel}=\lambda,\lambda_{\perp}=\lambda^{-0.5}$, as discussed in the previous section.
@@ 7 Discussion
To tackle this problem, we first analyze the fundamental condition of metric compatibility between adjacent director fields. We then study the compatible director fields for a general $n$-fold vertex, characterized into the generic case and the degenerate case. We prove that for the generic case there are two and only two compatible director fields which are orthogonal to each other (orthogonal duals), while for the degenerate case, there is a continuous family of compatible director fields. The Gaussian curvature of the actuated vertex is also computed. We develop a rigorous continuum mechanics framework to study the kinematics of isolated three-fold and four-fold vertices. Building on this vertex-level understanding, we construct two classes of large-scale director patterns that are compatible both before and after actuation. In other words, the directors satisfy metric compatibility at each crease, and the resulting global deformation is geometrically compatible. Although the general compatibility problem remains open (for example, a three-fold compatible director pattern that develops negative Gaussian curvature is always incompatible in the actuated configuration), our two proposed designs demonstrate broad and flexible design spaces. The first is a quadrilateral director pattern with free directors on the two boundaries. Using a marching algorithm and a rigidity theorem, the director information propagates to interior nodes, producing a compatible actuated configuration characterized by alternating positive and negative Gaussian curvature at the vertices. The second design combines three-fold and four-fold intersections, leveraging a folding angle theorem that shows the folding angle at a three-fold node depends solely on its adjacent directors. In this configuration, positive Gaussian curvature arises at the three-fold vertices and negative curvature at the four-fold vertices after actuation. On the application side, we propose a design strategy for active metamaterials based on periodic quadrilateral director patterns, which reveals two modes of motion by folding and stimulating.
@@ Acknowledgments
This work is based on discussions with Prof. Paul Plucinsky at USC, while we were working together at the University of Minnesota. We also thank Mr. Xi SU from PKU for providing the Householder matrix approach for the proof of Theorm 3.1. F.F. acknowledges the financial support from the National Natural Science Foundation of China (Grant No. 12472061).